# A Comprehensive Evaluation of Semantic Relation Knowledge of Pretrained Language Models and Humans

### v1

@@ 4.1 Prompt Design
Note that some prompts do not require any determiner, but others do. The notation [det] expresses that either “an” or “a” is chosen, based on certain conditions to be discussed later. Our prompts are formulated such that [w] always preceeds [v], and that there is no token after the [v] slot.
@@ 4.2 Target Words and Probes
In order to obtain our set of tuples $T^{r}$, we use existing word-relation-relatum tuples from Hyperlex [62] and from the category norm corpus by Overschelde et al [46]. We use only those tuples where both $w$ and $v$ are nouns, and where both are contained in the intersection of the vocabularies of all models that will be tested in the experiment. This resulted in a total of 1,347 tuples: 147 for hypernymy, 805 for hyponymy, 234 for meronymy, 52 for antonymy and 109 for synonymy. Note that none of the sources contributed any holonymy tuples.
@@
Symmetric augmentation can also be applied to those relations that have a reverse relation. If $(r_{1},r_{2})$ is a pair of reverse relations, the following holds:
@@
Hypernymy and hyponymy form a reverse relation pair; holonymy and meronymy form another reverse relation pair. A small change to the symmetric augmentation procedure is necessary. For relations $r_{1},r_{2}$ in a reverse relation pair, symmetric augmentation proceeds as follows:
@@
For example, we can reverse the meronymy tuple (“building”, MER, “wall”) to obtain a new holonymy tuple (“wall”, HOL, “building”). Note that if there are any duplicate tuples, they are removed to form the set $T^{r,aug}$. We will simplify notation after augmentation and use $T^{r}$ to refer to $T^{{r},aug}$. After symmetric augmentation, the total number of tuples has risen 1.66 fold (2,242, from 1,347). For holonomy, this process creates the only tuples in existence (186 tuples).
@@
Target words can now be extracted from tuples in the obvious way. For each relation $r$, we form $W^{r}$ as the set of target words $w$ from all tuples in $T^{r}$. Except for duplicates, each tuple contributes a target word. $W$, the union of $W^{r}$ for different relations $r$, denotes the set of unique target words in all experiments, independent of relation.
@@
The total number of probes we create is 10,546; for each relation, the component is the product of prompts and target words. Statistics for each relation can be gleaned from Table 1999The total over target words reported in the table is the sum over $|W^{r}|$. Because some target words are associated with more than one relation, this sum is different from $W$, the total number of unique target words, which is 1,266..
@@
Relation $|P^{r}|$ (prompts) $|W^{r}|$ (target words) $|X^{r}|$ (probes)
Hypernymy (HYP) 7 692 4,844
Hyponymy (HPO) 4 310 1,240
@@ 4.3 Relatum Sets
So far, we have constructed probes that we will give as inputs to models and as stimuli to humans. We now want to create $r$-relatum sets $Y^{r}$ that we can use for evaluation, for each target word $w\in W$ and relation $r$. We start by considering which properties good gold standards for our task would have.
@@
Unfortunately, the current relatum sets do not fulfill these two criteria. On average, they contain only 1.2 (antonymy) to 2.6 (hyponymy) relata per target word, and the average number of relations associated with each target word is only 1.3. This means that a majority of target words is associated with only one relation, making it impossible to assess models’ ability to distinguish between relations. Therefore, it is desirable to increase the number of relata per relatum set, as well as the number of associated relations for each target word.
For a given relation $r$ and a target word $w$, we increase the $r$-relata in the $r$-relatum set $Y^{r}$ as follows. We first retrieve all possible word senses of $w$ in WordNet. Then, for each word sense that has $r$-relata documented in WordNet, we update $Y^{r}$ by adding the new $r$-relatum, unless it is not included in the models’ shared vocabulary. In other words, all possible $r$-relata of any sense of the target word are included in the expanded $Y^{r}$. For example, “ending/1” is a synonym of “termination/4”, and “ending/3” is a synonym of “conclusion/6”. The resulting synonym set for ending therefore includes both “termination” and “conclusion”. This is so despite the fact that they refer to different senses of “ending” [^10] Note that the construction of synonym set for “termination” or “conclusion” results in a different expanded synonym set.. We additionally include indirect hypernyms and hyponyms of either of the target words senses, defined as those which lie within a path length of two in the WordNet hierarchy.
This procedure can result in a situation where more than one semantic relation holds between two word forms. For example, WordNet lists “conclusion/3” as a hyponym of “ending/3” and at the same time lists “conclusion/4” as a synonym of “ending/4”. The result are two relations holding between the word forms “conclusion” and “ending” [^11] The problem arises because our evaluation is performed at the word form level (as is the common approach), and not the sense level. If we were able to evaluate with senses disambiguated, we would be able to leave these relata in, with added profit.. We therefore solve this problem by removing all relationally ambiguous relata for each target word. After this step, we have a guarantee that for each target word, the relatum sets of different relations are mutually exclusive.
@@
Hypernymy (HYP) 1.2 $\pm$ 0.5 8.4 $\pm$ 6.8
Hyponymy (HPO) 2.6 $\pm$ 4.0 39.2 $\pm$ 54.9
Holonymy (HOL) 1.3 $\pm$ 0.6 2.9 $\pm$ 2.3
Meronymy (MER) 1.7 $\pm$ 3.4 3.9 $\pm$ 6.2
Antonymy (ANT) 1.1 $\pm$ 0.3 1.2 $\pm$ 0.5
Synonymy (SYN) 1.1 $\pm$ 0.2 3.5 $\pm$ 2.8
The expansion results in an increase in the average relatum set sizes, as can be seen from Table 2. The final average relatum set sizes range from 1.2 for antonymy to 39.2 for hyponymy. The average size of expanded hyponym sets is far larger than others because of the nature of hyponymy; as we descend the WordNet hierarchy to retrieve hyponyms, the number of hyponyms increases. The expansion also increases the average number of relations associated with each target word from 1.3 to 3.4.
These relatum sets constitute our gold standard in the upcoming evaluation. Of course it is possible for both humans and models to respond with a word that is not in the $r$-relatum set for either relation $r$. We call such words OOR (out of relatum set).
@@ 5 Metrics
The proposed evaluation framework consists of five metrics, two of which are novel. The novel metrics are called prototypicality and distinguishability. Prototypicality evaluates a property of semantic relations. Distinguishability evaluates agents’ ability to distinguish relations from each other. Soundness and completeness measure the performance of relatum prediction under the multiple relata setting. Symmetry and asymmetry have been studied before, but only with factual relations.
To calculate all metrics, we need a ranked list, for humans and for each model. The process starts with a probe, which we gain from target word $w^{r}$ and prompt $p^{r}$ by the verbalization function $\nu$. Using the probe, we elicit relata $v$ from multiple human participants, or from each model.
@@
From each $D(w^{r},p^{r};m)$ for either agent $m$, we can create a ranked list $L(w^{r},p^{r};m)$. The rank is established by the probability of that word from our distribution over relata $D(w^{r},p^{r};m)$. The list returned by models is as long as their vocabulary, so we need to introduce a cutoff $k$, which will be established separately for each metric. We denote $L_{k}(w^{r},p^{r};m)$ as the resulting response list for either agent $m$ is denoted. This allows us to treat human responses and model responses in a comparable way.
@@ 5.1 Soundness and Completeness
Note that we first average $\mathrm{Precision@1}$ scores over different prompts for the same target word. We then average over target words. Others before us [16, 50, 21] have used accuracy, which is mathematically identical to $\mathrm{Precision@1}$, but their numerical values are not comparable to our soundness. This is because we use relatum sets instead of single gold standard items. As any item of our relatum set counts as a hit, soundness values will be generally higher than the accuracy values in previous work.
@@
Completeness averages over the well-known information retrieval metric $\mathrm{Recall@k}$. Here, we set $k$ to the size of the relatum set or the size of the response list, whichever is smaller. Soundness and completeness values become identical when a relatum set only has one relatum.
5.2 Symmetry and Asymmetry
We measure whether reciprocal elicitation and forward elicitation happens for triplet $(w,r,v)$ using metrics proposed by Mruthyunjaya et al [42] (Equations (1) and (2), respectively). Our metrics differ from Mruthyunjaya et al’s in how averaging takes place. We average as we do in our calculation of the $\mathcal{S}$ and $\mathcal{C}$ scores. The summary statistic symmetry $\mathcal{M}_{k}(r;m)$ is achieved by first obtaining a symmetry score for each triplet, agent $m$, and relation $r$, and then averaging over the triplets.
$\mathcal{M}_{k}(w,r,v;m)$ $=\frac{1}{\lvert P^{r}\rvert}\sum_{p^{r}\in P^{r}}\mu_{k}(w,v,p^{r};m)$ (15)
$\mathcal{M}_{k}(r;m)$ $=\frac{1}{\lvert T^{r}\rvert}\sum_{(w,r,v)\in T^{r}}\mathcal{M}_{k}(w,r,v;m)$ (16)
where $r$ is either antonymy or synonymy and $\mathcal{M}_{k}(w,r,v;m)$ is the symmetry score for triplet $(w,r,v)$.
Asymmetry $\mathcal{A}_{k}(r;m)$ of agent $m$ on relation $r$ is calculated similarly as $\mathcal{M}_{k}(r;m)$.
$\mathcal{A}_{k}(w,r,v;m)$ $=\frac{1}{\lvert P^{r}\rvert}\sum_{p^{r}\in P^{r}}\alpha_{k}(w,v,p^{r};m)$ (17)
$\mathcal{A}_{k}(r;m)$ $=\frac{1}{\lvert T^{r}\rvert}\sum_{(w,r,v)\in T^{r}}\mathcal{A}_{k}(w,r,v;m)$ (18)
where $r$ is either hypernymy, hyponymy, holonymy, or meronymy.
For symmetric relations, the data at hand allows us to directly measure symmetry scores, for both models and humans. For asymmetric relations, we want to know if models successfully show only forward elicitation, and no backward elicitation on the same tuple. Note that for this we need a new probe, a trick probe, which expresses the opposite relationship to the original probe: on the same prompt template, it presents the relatum and then records whether backward elicitation happens, which it shouldn’t. For example, for the holonymy tuple $t^{\mathrm{HOL}}=$(“wall”, HOL, “building”), we need to create the trick probe “a building is a part of [det] [v] (in addition to what we already have, namely “a wall is a part of [det] [v]).
Trick probes can be constructed on the fly: the prompt from the original probe is used, and the target word is replaced with the relatum. This procedure yields 310$\times$7 trick probes for hypernymy, 692$\times$4 for hyponymy, 144$\times$7 for holonymy, and 186$\times$6 for meronymy.
On asymmetry, we evaluate only models. We assume that it is not necessary to evaluate humans, as they should have strong enough intuitions about the asymmetry of asymmetric relations.
@@ 5.3 Prototypicality
We now introduce metrics in order to determine the degree to which prototypicality is observed in the human responses. Note that unlike in our other experiments, we obtain human experimental performance that establishes the gold standard, rather than some pre-existing lexical data.
@@ 5.3.1 Response Entropy
$\mathcal{R}(w^{r},p^{r})=\begin{cases}-\sum_{v\in D(w^{r},p^{r};h)}\Pr(v)\frac{\log_{2}\Pr(v)}{\log_{2}\lvert D(w^{r},p^{r};h)\rvert},&\text{otherwise}\\ 0.&\text{if }\lvert D(w^{r},p^{r};h)\rvert=1\end{cases}$ (19)
As for all entropy-based metrics, lower numbers correspond to a stronger prototypicality effect. Maximal response entropy corresponds to a situation where all participants reply with the same single word, and nothing else. We define $\mathcal{R}(w^{r},p^{r})=0$ for this case [^13] The formula cannot be used in this case, as $\lvert D(w^{r},p^{r};h)\rvert=1$ and therefore $\log_{2}\lvert D(w^{r},p^{r};h)\rvert=0$.. Therefore, $\mathcal{R}(w^{r},p^{r})$ has a range between zero and one.
@@ 5.3.2 Prototypicality Score
We evaluate prototypicality of the model response by comparing the response with the human gold standard from above. In order to realize this, we need a similarity score that rewards models for satisfying the following requirements: 1) the prototype of the human response, i.e., the word most frequently elicited, is ranked highest in the model’s response, and 2) additionally, in the model’s response there are as many other words elicited from humans as possible, with similar rankings.
For example, consider the hypernymy probe “a wall is a part of [det] [v]. The human response is [“building”, “home”, “house”, “room”, …]. Given this ranked list, any model that returns the prototypical holonym “building” at the top position fulfills the first requirement. Concerning the other requirement, [“room”, “building”, “home”, “house”] is preferable to [“room”, “house”, “home”, “building”] because it preserves the precedence of “building” over “home” and “house” in the human response.
The first requirement can be implemented using the indicator function $\mathbb{I}[P]$. The second requirement can be implemented with the edit similarity $E(a,b)$ between word sequences $a$ and $b$, which is based on the edit distance [^14] The version of edit distance that we use here operates with insertion, deletion and substitution, with a weight of two for substitution and a weight of one for the other operations. The maximum value of edit distance is the weight of substitution times the length of the larger of the two sequences compared, which is bounded by $2k$. We therefore normalize the score by $2k$ and turn it into a similarity metric by reporting the distance from 1.. This time, $k$ differs for each probe; it is set to the number of words in human response for the probe in evaluation.
@@ 5.4 Distinguishability
If a model distinguishes relation $r$ well from relation $s$, then the ranks of $r$-relata in the response should be overall much lower than the ranks of $s$-relata. This takes into account an aspect that soundness does not. Consider the example in Figure 1. There are two responses, A and B, to the holonymy probe “A wall is a part of [det] [v].
@@
For an ordered pair of semantic relations $(r,s)$, we define $\delta(w^{r},p^{r},s;m)$ as mean relative rank of $s$-relata in the response to a probe of relation $r$. In addition, $\delta(r,s;m)$ is defined as the average of $\delta(w^{r},p^{r},s;m)$ over all prompts and target words. The distinguishability of $r$ from $s$, denoted by $\mathcal{D}(r,s;m)$, is the difference between $\delta(r,s;m)$ and $\delta(r,r;m)$ as follows.
@@
where $\rho(a,b)$ is the normalized relative rank of a word $a$ in a list $b$. Normalization of $\rho(a,b)$ by $k$ (here set to the size of the relatum set of target word $w$) results in a range $[0,1]$.
Higher $\mathcal{D}$ scores indicate better distinction. Note that this metric can incur negative values, namely if the highest-ranked correct relatum is ranked after an incorrect relatum. In that case, the model has commited an error so grave that we are no longer interested in the rest of the response. We therefore assign zero to all cases of negative difference.

### v4

@@ 4.1 Prompt Design
Note that some prompts do not require any determiner, but others do. The notation [det] expresses that either “an” or “a” is chosen, based on certain conditions to be discussed later. Our prompts are formulated such that [w] always precedes [v], and that there is no token after the [v] slot.
@@ 4.2 Target Words and Probes
In order to obtain our set of tuples $T^{r}$, we use existing word-relation-relatum tuples from Hyperlex [67] and the category norm corpus by Overschelde et al [49]. We use only those tuples where both $w$ and $v$ are nouns, and where both are contained in the intersection of the vocabularies of all models that will be tested in the experiment. This resulted in a total of 1,340 tuples: 145 for hypernymy, 800 for hyponymy, 234 for meronymy, 52 for antonymy and 109 for synonymy. Note that none of the sources contributed any holonymy tuples.
@@
Symmetric augmentation can also be applied to those relations that have a reverse relation. If $r_{1}$ and $r_{2}$ are reverse relations of each other, the following holds:
@@
Hypernymy and hyponymy are reverse relations of each other; so are holonymy and meronymy. A small change to the symmetric augmentation procedure is necessary. For reverse relations $r_{1}$ and $r_{2}$, symmetric augmentation proceeds as follows:
@@
For example, we can reverse the meronymy tuple (“building”, MER, “wall”) to obtain a new holonymy tuple (“wall”, HOL, “building”). Note that if there are any duplicate tuples, they are removed to form the set $T^{r,aug}$. We will simplify notation after augmentation and use $T^{r}$ to refer to $T^{{r},aug}$. After symmetric augmentation, the total number of tuples has risen 1.78 fold (2,390, from 1,340). For holonomy, this process creates the only tuples in existence (234 tuples). For other relations, the number of added tuples is 658 for hypernymy, 109 or synonymy, 46 for antonymy and 3 for hyponymy.
@@
Target words can be extracted from tuples as follows. For each relation $r$, we form $W^{r}$ as the set of target words $w$ from all tuples in $T^{r}$. Except for duplicates, each tuple contributes a target word. $W$, the union of $W^{r}$ for different relations $r$, denotes the set of unique target words in all experiments, independent of relation.
@@
The total number of probes we create is 10,507; for each relation, the component is the product of prompts and target words. Statistics for each relation can be gleaned from Table 1888The total over target words reported in the table is the sum over $|W^{r}|$. Because some target words are associated with more than one relation, this sum is different from $W$, the total number of unique target words, which is 1,266..
@@
Relation
$|P^{r}|$
$|W^{r}|$
(target words)
$|X^{r}|$
Hypernymy (HYP) 7 687 4,809
Hyponymy (HPO) 4 309 1,236
@@ 4.3 Relatum Sets
So far, we have constructed probes that we will give as inputs to models and as stimuli to humans. We now create $r$-relatum sets $Y^{r}$ that we can use for evaluation, for each target word $w\in W$ and relation $r$. We start by considering which properties good gold standards for our task would have.
@@
Unfortunately, the current relatum sets do not fulfill these two criteria. On average, they contain only 1.1 (antonymy) to 2.6 (hyponymy) relata per target word, and the average number of relations associated with each target word is only 1.5. This means that a majority of target words are associated with only one relation, making it impossible to assess models’ ability to distinguish between relations. Therefore, it is desirable to increase the number of relata per relatum set, as well as the number of associated relations for each target word.
For a given relation $r$ and a target word $w$, we increase the $r$-relata in the $r$-relatum set $Y^{r}$ as follows. We first retrieve all possible word senses of $w$ in WordNet. Then, for each word sense that has $r$-relata documented in WordNet, we update $Y^{r}$ by adding the new $r$-relatum, unless it is not included in the models’ shared vocabulary. In other words, all possible $r$-relata of any sense of the target word are included in the expanded $Y^{r}$. For example, “ending/1” is a synonym of “termination/4”, and “ending/3” is a synonym of “conclusion/6”. The resulting synonym set for ending therefore includes both “termination” and “conclusion”. This is so despite the fact that they refer to different senses of “ending” [^9] Note that the construction of synonym set for “termination” or “conclusion” results in a different expanded synonym set.. We additionally include indirect hypernyms and hyponyms of either of the target word senses, defined as those which lie within a path length of two in the WordNet hierarchy.
This procedure can result in a situation where more than one semantic relation holds between two word forms. For example, WordNet lists “conclusion/3” as a hyponym of “ending/3” and at the same time lists “conclusion/4” as a synonym of “ending/4”. The result is two relations holding between the word forms “conclusion” and “ending” [^10] The problem arises because our evaluation is performed at the word form level (as is the common approach), and not the sense level. If we were able to evaluate with senses disambiguated, we would be able to leave these relata in, with added profit.. We therefore solve this problem by removing all relationally ambiguous relata for each target word. After this step, we have a guarantee that for each target word, the relatum sets of different relations are mutually exclusive.
@@
Hypernymy (HYP) 1.2 $\pm$ 0.5 8.9 $\pm$ 7.1
Hyponymy (HPO) 2.6 $\pm$ 4.0 20.5 $\pm$ 43.0
Holonymy (HOL) 1.3 $\pm$ 0.6 2.5 $\pm$ 2.1
Meronymy (MER) 1.7 $\pm$ 3.5 2.9 $\pm$ 4.4
Antonymy (ANT) 1.1 $\pm$ 0.3 1.1 $\pm$ 0.4
Synonymy (SYN) 1.1 $\pm$ 0.2 3.0 $\pm$ 2.7
The expansion increases the average relatum set sizes, as can be seen from Table 2. The final average relatum set sizes range from 1.1 for antonymy to 20.5 for hyponymy. The average size of expanded hyponym sets is far larger than others because of the nature of hyponymy; as we descend the WordNet hierarchy to retrieve hyponyms, the number of hyponyms increases. The expansion also increases the average number of relations associated with each target word from 1.1 to 2.9.
These relatum sets constitute our gold standard in the upcoming evaluation. Of course, both humans and models can respond with a word that is not in the $r$-relatum set for either relation $r$. We call such words OOR (out of the relatum set).
@@ 5 Metrics
The proposed evaluation framework consists of five metrics, two of which are novel. The novel metrics are called prototypicality and distinguishability. Prototypicality evaluates a property of semantic relations. Distinguishability evaluates agents’ ability to distinguish relations from each other. Soundness and completeness measure the performance of relatum prediction under the multiple relata setting. Symmetry has been studied before, but only with factual relations.
To calculate all metrics, we need a ranked list, for humans and each model. The process starts with a probe, which we gain from the target word $w^{r}$ and prompt $p^{r}$ by the verbalization function $\nu$. Using the probe, we elicit relata $v$ from multiple human participants, or from each model.
@@
From each $D(w^{r},p^{r};m)$ for either agent $m$, we can create a ranked list $L(w^{r},p^{r};m)$. The rank is established by the probability of that word from our distribution over relata $D(w^{r},p^{r};m)$. The list returned by models is as long as their vocabulary, so we need to introduce a cutoff $k$, which will be established separately for each metric. We denote $L_{k}(w^{r},p^{r};m)$ as the resulting response list with the top $k$ words for the agent $m$. This allows us to treat human responses and model responses in a comparable way.
@@ 5.1 Soundness and Completeness
Note that we first average $\mathrm{Precision@1}$ scores over different prompts for the same target word. We then average over the target words. Previous research [20, 53, 24] used accuracy, which is mathematically identical to $\mathrm{Precision@1}$, but their numerical values are not comparable to our soundness. This is because we use relatum sets instead of single gold standard items. As any item of our relatum set counts as a hit, soundness values will be generally higher than the accuracy values in the previous work.
@@
Completeness averages over the well-known information retrieval metric $\mathrm{Recall@k}$. Here, we set $k$ to the size of the relatum set or the size of the response list, whichever is smaller. Soundness and completeness values become identical when the relatum set is singleton.
5.2 Symmetry
We measure whether reciprocal elicitation happens for tuple $(w,r,v)$ using metrics proposed by Mruthyunjaya et al [45] (Equation (1)). Our metrics differ from Mruthyunjaya et al’s in how averaging takes place. We average as we do in our calculation of the $\mathcal{S}$ and $\mathcal{C}$ scores. The summary statistic symmetry $\mathcal{M}_{k}(r;m)$ is achieved by first obtaining a symmetry score for each tuple, agent $m$, and relation $r$, and then averaging over the tuples.
$\mathcal{M}_{k}(w,r,v;m)$ $=\frac{1}{\lvert P^{r}\rvert}\sum_{p^{r}\in P^{r}}\mu_{k}(w,v,p^{r};m),$ (14)
$\mathcal{M}_{k}(r;m)$ $=\frac{1}{\lvert T^{r}\rvert}\sum_{(w,r,v)\in T^{r}}\mathcal{M}_{k}(w,r,v;m),$ (15)
[…]
@@ 5.3 Prototypicality
We now introduce metrics in order to determine the degree to which prototypicality is observed in the human responses. Note that unlike in our other metrics, we obtain human experimental performance that establishes the gold standard, rather than some pre-existing lexical data.
@@ 5.3.1 Response Entropy
[…]
As for all entropy-based metrics, lower numbers correspond to a stronger prototypicality effect. Maximal response entropy corresponds to a situation where all participants reply with the same single word, and nothing else. We define $\mathcal{R}(w^{r},p^{r})=0$ for this case. Therefore, $\mathcal{R}(w^{r},p^{r})$ has a range between zero and one.
@@ 5.3.2 Prototypicality Score
We evaluate the prototypicality of the model response by comparing it with the human gold standard, which we will establish experimentally later. In order to realize this, we need a similarity score that rewards models for satisfying the following requirements: 1) the prototype of the human response, i.e., the word most frequently elicited, is ranked highest in the model’s response, and 2) additionally, in the model’s response there are as many other words elicited from humans as possible, with similar rankings.
For example, consider the hypernymy probe “a wall is a part of [det] [v]”. The human response is [“building”, “home”, “house”, “room”, …]. Given this ranked list, any model that returns the prototypical holonym “building” at the top position fulfills the first requirement. Concerning the other requirement, [“room”, “building”, “home”, “house”] is preferable to [“room”, “house”, “home”, “building”] because it preserves the precedence of “building” over “home” and “house” in the human response.
[…]
@@ 5.4 Distinguishability
If a model distinguishes relation $r$ well from relation $s$, then the ranks of $r$-relata in the response should be overall much lower than the ranks of $s$-relata.
Consider the example in Figure 1. There are two responses, A and B, to the holonymy probe “A wall is a part of [det] [v]”.
@@
For an ordered pair of semantic relations $(r,s)$, we define $\delta(w^{r},p^{r},s;m)$ as a mean relative rank of $s$-relata in the response to a probe of relation $r$. In addition, $\delta(r,s;m)$ is defined as the average of $\delta(w^{r},p^{r},s;m)$ over all prompts and target words. The distinguishability of $r$ from $s$, denoted by $\mathcal{D}(r,s;m)$, is the difference between $\delta(r,s;m)$ and $\delta(r,r;m)$ as follows.
@@
where $\rho(a,b)$ is the normalized relative rank of a word $a$ in a list $b$. Normalization of $\rho(a,b)$ by $k$ (here set to the size of the relatum set of target word $w$) results in a range [0,1][0,1][ 0 , 1 ].
Higher $\mathcal{D}$ scores indicate better distinction. Note that this metric can incur negative values, namely if the highest-ranked correct relatum is ranked after an incorrect relatum. In that case, the model has committed an error so grave that we are no longer interested in the rest of the response. We therefore assign zero to all cases of negative difference.